OpenStudy (anonymous):

(y^2-81/6y+12)*(y+2/y-9)

jimthompson5910 (jim_thompson5910):

Hints: y^2 - 81 factors to (y-9)(y+9) 6y+12 factors to 6(y+2)

OpenStudy (luigi0210):

do you understand what he means?

OpenStudy (anonymous):

I think so. It would become (y-9)/6, right?

OpenStudy (whpalmer4):

Close, but not quite right....

OpenStudy (luigi0210):

try again kiddo :P

OpenStudy (whpalmer4):

\[\frac{(y^2-81)}{(6y+12)}*\frac{(y+2)}{(y-9)} = \frac{(y-9)(y+9)}{6(y+2)}*\frac{(y+2)}{(y-9)} =\]